OpenStudy (anonymous):

if logb^2 = x and logb^3 = y, evaluate x and y logb432 logb648 logb 4/?9 logb81/logb^2

OpenStudy (anonymous):

is this \[\log_b(2)=x,\log_b(3)=y\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use \[432=2^43^3\] to get \[\log_b(432)=\log_b(2^4\times3^3)=\log_b(2^4)+\log_b(3^3)=4\log_b(2)+3\log_b(3)\] \[4x+3y\] a problem only a math teacher could love. others are similar

OpenStudy (anonymous):

by which i mean write the input in terms of powers and multiples of 2 and 3, then break the log apart as above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me out with the rest especially logb 4/9 and logb8^1/logb^2

OpenStudy (anonymous):

\[\frac{4}{9}=\frac{2^2}{3^2}\] so \[\log(\frac{4}{9})=\log(\frac{2^2}{3^2})=\log(2^2)-\log(3^2)=2\log(2)-2\log(3)=2x-2y\]

OpenStudy (anonymous):

ok